120
If the two events are independent then the probability of them both happening is Pr(event1) X Pr(event2).
Which in your case is 0.75x0.50=0.375 which translates into 37.5%
Rides a bike to school!
Rides a bike to school!
fdswewqa
What is the probability that all four 1 seeded teams make it to the Final Four?
It happened for the first time in 2008. To calculate the probability you'd need to know the probability of each number 1 seed winning each game, and that changes every year (and can't really be measured).
Define the probability of something happening?
The probability that something will happen is the number of times that that something can occur divided by the number of times that anything can occur.
An example would be to consider a standard deck of 52 cards. There are 4 sevens in the deck, so the probability of drawing a seven is 4 in 52, or 1 in 13, or about 0.0769.
people eat chees often
well if you take the four out of the seven its 4/7 convert that into a percent wich equals 57%
You should ask the dealer what the proper indication is. Usually the desire to be hit in Blackjack is indicated by bending a finger towards you, in a "come here" sign, or by pointing down at the cards, but this could be confused with the stay command which is usually a flat open hand over the cards. Go with my first recommendation and ask the dealer.
It is a progression of terms whose reciprocals form an arithmetic progression.
1/13
What is expected level of sales if there is a 15 probability the economy will be strong?
The answer will depend on the product. If the economy is strong the sales of inferior goods will fall. Also, it depends on the level at which the sales are currently.
Three out of twelve
8 were defective while (75-8=67) weren't. So the estimated probability that a flyer won't be defective is 67/75
What is the probability of getting 2 tails if 3 fair coins are tossed?
There are eight possible outcomes: HHH, HHT, HTT, HTH, TTT, TTH, THH, THT. Of these, 3 contain two tails: HTT, TTH, and THT and the probability of getting two tails is 3/8. If the question were 'getting at least two tails' then TTT would need to be included for a probability of 4/8 or 0.5.
If 3 fair coins are tossed what is the probability of getting exactly 2 tails?
The probability is 3/8 = 0.375
Let's break this down and see what we can do with it. First, the probability of transition of an atom from the initial state (m) to final state (n) will depend on what it is that is transitioning, and also on how it's transitioning. Which element - and which isotope - is transitioning? And how it is transitioning? Under what conditions? Radioactive decay? Neutron (or other particle) absorption? With the stating atom and the mode of change, we can determine the probability of transition as was asked. With the first part of the information in hand, we can determine the energy loss in the heavy charged particle emission event. We should also be able to determine the time over which the event takes place, and then discover the rate of energy loss (though it will represent only an average) for said event.
What is the range of a number?
The range of a number set is the difference between the highest number and the lowest number in the set. To find the range, you subtract the smallest number from the largest number.
What is the probability of getting dealt two aces in a deck of cards?
The odds are 220:1 of being dealt pocket aces.
If four fair coins are tossed what is the probability of getting at least two heads?
1 - P (1 or 0) =
1 - P(1) - P(0) =
1 - ((4 above 1) * 0.5 * 0.5^3) - ((4 above 0) *0.5^4) =
1 - (4/16) - (1/16) =
1- (5/16) =
11/16
zero...
Whats the probability of getting at most one heads when two fair coins are tossed?
When two fair coins are tossed, you have the following possible outcomes:
HH, HT, TH, TT.
So, at most implies that you get either i) zero heads or ii) one head.
From the possible outcomes we see that 3 times we satisify the outcome.
Thus, probability of at most one head is 3/4.
Probability of a coin landing on heads if done 100 times?
The probability is 0.5 regardless how many times you toss the coin."
To get the answer is simple. You just convert 67% to a decimal. It is 0.67. Then multiply by 5
0.67x 0.67 x 0.67 x 0.67 x 0.67 = 0.135